OpenStudy (anonymous):

1 over 4x-1=7 whats the solution to the equation ?

jhonyy9 (jhonyy9):

1 ---- =7 4x-1 - is correct ?

jhonyy9 (jhonyy9):

common denominator is 4x-1 than multiplie 7 by it 1 =7(4x-1) 1=28x-7 8=28x x=8/28 x=2/7

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

i guess .. thanks

jhonyy9 (jhonyy9):

yw good luck bye